# Probabilistic Guarantees to Explicit Constructions: Local Properties of Linear Codes

### v1

@@ Abstract
We present a general framework for derandomizing random linear codes with respect to a broad class of permutation-invariant properties, known as local properties, which encompass several standard notions such as distance, list-decoding, list-recovery, and perfect hashing. Our approach extends the classical Alon–Edmonds–Luby (AEL) construction through a modified formalism of local coordinate-wise linear (LCL) properties, introduced by Levi, Mosheiff, and Shagrithaya (2025). The main theorem demonstrates that if random linear codes satisfy the complement of an LCL property $\mathcal{P}$ with high probability, then one can construct explicit codes satisfying the complement of $\mathcal{P}$ as well, with an enlarged yet constant alphabet size. This gives the first explicit constructions for list recovery, as well as special cases (e.g., list recovery with erasures, zero-error list recovery, perfect hash matrices), with parameters matching those of random linear codes. More broadly, our constructions realize the full range of parameters associated with these properties at the same level of optimality as in the random setting, thereby offering a systematic pathway from probabilistic guarantees to explicit codes that attain them. Furthermore, our derandomization of random linear codes also admits efficient (list) decoding via recently developed expander-based decoders.
@@ 1.1 Our Results
For any reasonable LCL property $\mathcal{P}$, there exists a suitable (linear) inner code, a bipartite expander, and an outer code such that the AEL procedure, when instantiated with these components, yields an explicit linear code $\mathcal{C}_{\textrm{AEL}}$ that does not satisfy $\mathcal{P}$, and whose rate is arbitrarily close to the threshold rate.
We observe that our explicit codes achieve parity with random linear codes in terms of all parameters associated with the LCL property.
@@
Two important tradeoffs arise in our construction: alphabet size and the underlying field of linearity. First, our construction incurs an exponential increase in alphabet size. This phenomenon is characteristic of all constructions based on the AEL procedure in the literature, and our setting is no exception. Second, while our codes are defined over a larger field, they remain linear only with respect to a subfield—namely, the field over which the inner code is defined. This limitation, however, does not pose difficulties for most applications.
@@
[LMS25] also proved that random Reed-Solomon codes and random linear codes are equivalent with respect to local properties: that is, they have the same threshold rates for all reasonable local properties. This implies that our constructions also match the parameters attained by random Reed-Solomon codes for reasonable local properties, with the additional property of having constant alphabet size.
@@
List-decodable and list-recoverable codes have found uses in numerous areas of theoretical computer science, including pseudorandomness [Tre99, GUV09, LP20], compressed sensing [NPR12], and algorithms [LNNT19, DW22]. Their broad applicability has motivated the development of several explicit constructions spanning a wide range of parameter regimes. For example, list-recoverable codes have been used in constructions of list-decodable codes [GI03, GR06b, KRSW23], and locally decodable codes [HRW20].
@@
In the case of list-recovery, the list-recovery capacity theorem (see [Res20], Proposition 2.4.14 for a proof) states that there exist codes that are $(\rho,\ell,L)$-list-recoverable, with
@@
A long line of works (e.g., [RW18], [LP20], [GLS+24], [LS25]) have studied list-recovery in the large alphabet regime. [GLS+24] showed that random Reed-Solomon codes are $(1-R-\varepsilon,\ell,O(\ell/\varepsilon))$-list recoverable codes with rate $\Omega(\varepsilon/(\sqrt{\ell}\log(1/\varepsilon)))$. In [LS25], the authors showed for any rate $R$, random linear codes are $(1-R-\varepsilon,\ell,L)$-list recoverable, where $L\leq\left(\frac{\ell}{\varepsilon}\right)^{O\left(\frac{\ell}{\varepsilon}\right)}$.
@@
For list-recoverability, [KRSW23] (Theorem 6.7), along with Tamo’s analysis of the output list size of Folded Reed-Solomon codes [Tam24] (Theorem 4.5) gives explicit constructions of codes over alphabet $\Sigma$ that are $(1-R-\varepsilon,\ell,L)$-list-recoverable, with $\left|\Sigma\right|\leq\ell^{O(1/\varepsilon^{4})}$, and $L\leq\left(\ell/\varepsilon\right)^{\left(\frac{\ell}{\varepsilon}\right)^{2}\cdot\log\left(\frac{\ell}{\varepsilon}\right)}$.[ST25] gives $(1-R-\varepsilon,\ell,L)$-list-recoverable codes, where $\left|\Sigma\right|=L\leq\exp\left((\ell/\varepsilon)^{O\left(\log(\ell/\varepsilon)\right)}\right)$. Our result gives constructions with smaller output list sizes than either result, at the cost of increased alphabet size. All results mentioned below follow from Theorem˜1.1 by specializing to the appropriate list recovery variant.
Theorem 1.4 (Informal, see Corollary˜6.10, Corollary˜6.9).
@@
Instantiating the codes with the best known upper bound on list sizes from [LS25], we get explicit $(1-R-\varepsilon,\ell,L)$-list recoverable codes with rate $R-2\varepsilon$, and $L\leq(\ell/\varepsilon)^{O\left(\ell/\varepsilon\right)}$, with alphabet size at most $\exp\bigl((\ell/\varepsilon)^{(\ell/\varepsilon)^{(\ell/\varepsilon)}}\bigr)$. The generality of our main result also yields explicit constructions for related notions such as zero-error list recovery and erasure list recovery. Zero-error list recovery is a special case of list recovery in which the decoding radius is zero. Such codes have found applications in the design of data structures for the heavy hitters problem [DW22]. In the case of erasure list recovery, some of the input lists may contain only the blank symbol, and the objective is to minimize the number of codewords that remain consistent with the non-blank input lists.
Theorem 1.5 (Informal, see Corollary˜6.11).
@@
Theorem 1.6 (Informal, see Corollary˜6.12).
@@
Theorem 1.7 (Informal, see Corollary˜6.13).
@@
Lastly, we get a modest improvement in the alphabet size of explicit list-decodable codes approaching the Generalized Singleton Bound. Previously, [JMST25] constructed explicit codes with the same parameters, except with alphabet size equal to $2^{{\mathrm{poly}}(L^{L}/\varepsilon)}$.
@@
The proof in [JMST25] is based on induction, and leverages the expander property of the graph used in the construction. In contrast, ours is a proof by contradiction, and relies on the sampling property of the underlying graph, thereby providing an alternative proof of essentially the same result. Although implied by Theorem˜1.1, we include a proof of Theorem˜1.8 as it more transparently illustrates the ideas behind the proof of the former theorem.
@@ 1.2 Organization
In Section˜2, we provide an overview of the proofs for the list-decoding case (Theorem˜1.8) and the general result (Theorem˜1.1). The detailed proofs appear in Section˜4 and Section˜5, respectively. Section˜5 is independent of Section˜4; therefore, readers interested solely in the general theorem’s proof may proceed directly to it. Finally, the results pertaining to list-recovery variants and perfect hash matrices (Theorem˜1.4, Theorem˜1.5, Theorem˜1.6, Theorem˜1.7) are presented in Section˜6.
@@ 2 Technical Overview
We begin with an overview of the proof of Theorem˜1.8, which addresses the special case of list decoding. We instantiate the AEL procedure with the following three components: an inner code of constant block length, found through brute force, an explicit outer code having high rate and distance $\delta_{\textrm{out}}$, where $\delta_{\textrm{out}}$ is a constant, and an explicit bipartite expander graph $G$. We note that $G$ can equivalently be interpreted as a sampler, a viewpoint that has been leveraged in prior works (cf. [KMRS17, KRSW23]) to analyze AEL-based constructions. This sampling perspective naturally motivates a strategy for the explicit construction of codes with strong list-decoding parameters: perform a brute-force search to identify a constant block-length code with good list-decoding parameters, and then use this code as the inner code in the AEL procedure to obtain $\mathcal{C}_{\textrm{AEL}}$. If $\mathcal{C}_{\textrm{AEL}}$ has pairwise distinct codewords $c_{1},\ldots,c_{L+1}$ close to some received word $y$, then the codewords agree with $y$ at a large number of coordinates. The nature of these agreements can be encoded by an agreement hypergraph on $L+1$ vertices that contains $n$ hyperedges, one for each coordinate. The hyperedge for a coordinate is simply the set of indices of codewords agreeing with $y$ on that coordinate. It follows that if $c_{1},\ldots,c_{L+1}$ agree with $y$ on many coordinates, then the sum of the sizes of the hyperedges is large. Let these agreements be described by an agreement hypergraph denoted by $\mathcal{H}$. Upon invoking the sampling property of the graph $G$, it is observed that the precise pattern of these agreements is “ported over” to a significant number of vertices on the left. As a result, the projections of codewords $c_{1},\ldots,c_{L+1}$ and $y$ onto several left vertices have agreements that closely resemble the agreement pattern described by $\mathcal{H}$. Consequently, the codeword projections agree with the projection of $y$ at a large number of coordinates within the inner code. But as the codeword projections are also codewords of the inner code, and because the inner code has good list-decoding parameters, its codewords do not have a lot of agreements with the local projection of $y$. Therefore, we have seemingly arrived at a contradiction.
@@
We now search for a linear code of constant block length satisfying the above property, and use it as an inner code. The analysis now proceeds in a manner analogous to the first approach. If $L+1$ codewords from $\mathcal{C}_{\textrm{AEL}}$ have a large number of agreements with some received word $y$, then the agreement hypergraph contains a weakly-partition-connected hypergraph $\mathcal{H}$. Consequently, a subset of those codewords satisfy the constraints as set forth by $\mathcal{H}$. We then use the sampling property of the graph $G$ to port over information about the nature of the agreements to the left side, with the result that the inner codewords at several vertices on the left satisfy the constraints described by $\mathcal{H}$. From the parameters that we eventually set, it is seen that the number of such parts exceeds $n(1-\delta_{\textrm{out}})$ (where $\delta_{\textrm{out}}$ is the minimum distance of the outer code). Consequently, at least one set of inner codewords from such a part must be non-trivial, thereby contradicting our assumption on the inner code.
@@
Every linear constraint set can be written down as a matrix of dimension $L\times L$. Two linear constraint sets $S$ and $T$ are said to be of the same type if there exists a full rank linear transformation that maps the matrix associated with $S$ to that associated with $T$. Even though the number of types can be exponential in $L^{2}$, it is independent of the block length of $\mathcal{C}_{\textrm{AEL}}$. This fact ensures that if there are $L$ pairwise distinct codewords in $\mathcal{C}_{\textrm{AEL}}$ that satisfy the linear constraints described by some local profile $M$ associated with $\mathcal{P}$, then upon arranging them in an $n\times L$ matrix, a significant fraction of the coordinates can be partitioned according to the constraint type satisfied by the row whose index is equal to the coordinate. Moreover, the fractional size of each of these sets will be a constant. Thus, the sampling property of the graph $G$ can be utilized to port over the proportion of these types onto several projections on the left. The local projections, which are also codewords belonging to the inner code, consequently have codewords that satisfy (a close approximation of) the local profile $M$. If the AEL procedure were instantiated with an inner code that avoids vector sets satisfying close approximations of the local profiles associated with $\mathcal{P}$, one might hope to derive a contradiction. However, we encounter the same obstacle as before: the guarantee for the inner code applies only when the inner codewords are pairwise distinct, and we lack a mechanism to ensure that this condition is met.
@@
Recall that we are interested in codes that satisfy the complement of an LCL property $\mathcal{P}$. We shall prove in Section˜5 that if random linear codes of rate $R$ satisfy the complement of $\mathcal{P}$, then random linear codes of rate $R$ also satisfy the following property: with high probability, they do not contain any non-zero matrices satisfying any implied local profile $I(M)$, where $M$ is any local profile associated with $\mathcal{P}$. Upon instantiating the AEL procedure with an inner code satisfying the aforementioned property, the analysis can be carried out in a manner analogous to that of the naive approach detailed above. If $L$ pairwise distinct codewords in $\mathcal{C}_{\textrm{AEL}}$ satisfy a local profile $M$ associated with $\mathcal{P}$, then upon arranging these codewords in a $n\times L$ matrix and applying the appropriate linear map to each of its rows, the matrix obtained satisfies the constraints set forth by the implied local profile $I(M)$. Upon porting the constraint types to the left as before, we see that the projected matrices on to several left vertices are codewords belonging to the inner code, and moreover, they satisfy I(M)I(M) [^2] In reality, the projected matrices satisfy a close approximation of $I(M)$, instead of $I(M)$. Nevertheless, the inner codes we employ are chosen to be sufficiently robust to accommodate this technical subtlety.. The parameters are instantiated in a way so as to ensure that this occurs on more than $(1-\delta_{\textrm{out}})n$ left vertices, and therefore there is one projected matrix that is non-zero. But because the columns of the projected matrices are codewords of the inner code, and because the inner code avoids containing all non-zero matrices satisfying $I(M)$ for any $M$ associated with the LCL property $\mathcal{P}$, we successfully arrive at a contradiction.

### v2

@@ Abstract
We present a general framework for derandomizing random linear codes with respect to a broad class of properties, known as local properties, which encompass several standard notions such as distance, list-decoding, list-recovery, and perfect hashing. Our approach extends the classical Alon–Edmonds–Luby (AEL) construction through a modified formalism of local coordinate-wise linear (LCL) properties, introduced by Levi, Mosheiff, and Shagrithaya (2025). The main theorem demonstrates that if random linear codes satisfy the complement of an LCL property $\mathcal{P}$ with high probability, then one can construct explicit codes satisfying the complement of $\mathcal{P}$ as well, with an enlarged yet constant alphabet size. This gives the first explicit constructions for list recovery, as well as special cases (e.g., list recovery with erasures, zero-error list recovery, perfect hash matrices), with parameters matching those of random linear codes. More broadly, our constructions realize the full range of parameters associated with these properties at the same level of optimality as in the random setting, thereby offering a systematic pathway from probabilistic guarantees to explicit codes that attain them. Furthermore, our derandomization of random linear codes also admits efficient (list) decoding via recently developed expander-based decoders.
@@ 1.1 Our Results
For any LCL property $\mathcal{P}$, there exists a suitable (linear) inner code, a bipartite expander, and an outer code such that the AEL procedure, when instantiated with these components, yields an explicit linear code $\mathcal{C}_{\textnormal{AEL}}$ that does not satisfy $\mathcal{P}$, and whose rate is arbitrarily close to the threshold rate.
A few remarks about the result are in order.
Remark 1.2 (Optimality, LDPC property).
We observe that our explicit codes achieve parity with random linear codes in terms of all parameters associated with the LCL property. Moreover, our explicit constructions are LDPC codes: meaning that each row of the parity check matrix only has constant number of non-zero entries. This property has previously been useful for designing very fast encoding and decoding algorithms.
Remark 1.3 (Combining LCL properties).
The nature of LCL properties enables one to “combine” a list of LCL properties $\mathcal{P}_{1},\mathcal{P}_{2},\ldots$ to create a single LCL property $\mathcal{P}$. Upon applying Theorem˜1.1 to $\mathcal{P}$, we obtain an explicit construction that does not satisfy $\mathcal{P}_{1},\mathcal{P}_{2},\ldots$ simultaneously. We refer the reader to Corollary˜6.2 for more details.
@@
Two important tradeoffs arise in our constructions: alphabet size and the underlying field of linearity.
    Increased alphabet size: Our constructions require an increased alphabet size, with the increase being exponential in ${q_{0}}^{{\mathrm{poly}}(L/\varepsilon)}$, where $q_{0}$ is the size of the field whose elements are used to specify the constraints of the property, $L$ is the locality, and $\varepsilon$ is the gap between the rate of the explicit construction and the threshold rate. This phenomenon is characteristic of several constructions based on the AEL procedure in the literature, and our setting is no exception.
    Linearity over a subfield: While our codes are defined over a larger field, they remain linear only with respect to a subfield—namely, the field over which the inner code is defined. This limitation, however, does not pose difficulties for most applications.
@@
[LMS25] also proved that random Reed-Solomon codes and random linear codes are equivalent with respect to LCL properties: that is, they have the same threshold rates for all LCL properties. This implies that our constructions also match the parameters attained by random Reed-Solomon codes for all LCL properties, with the additional property of having constant alphabet size.
@@
List-decodable and list-recoverable codes have found uses in numerous areas of theoretical computer science, including pseudorandomness [Tre99, GUV09, LP20], compressed sensing [NPR12], and algorithms [LNNT19, DW22]. Their broad applicability has motivated the development of several explicit constructions spanning a wide range of parameter regimes. For example, list-recoverable codes have been used in constructions of list-decodable codes [GI03, GR06b, KRSW23], and locally decodable codes [HRW20]. For a comprehensive overview of applications of list-recoverable codes, we refer the reader to the recent survey by Resch and Venkitesh [RV25].
@@
In the case of list-recovery, the list-recovery capacity theorem (see [Res20], Proposition 2.4.14 for a proof) establishes the existence of codes that are $(\rho,\ell,L)$-list-recoverable, with
@@
A long line of works (e.g., [RW18], [LP20], [GLS+24], [LS25]) have studied list-recovery in the large alphabet regime. [GLS+24] showed that random Reed-Solomon codes are $(1-R-\varepsilon,\ell,O(\ell/\varepsilon))$-list recoverable codes with rate $\Omega(\varepsilon/(\sqrt{\ell}\log(1/\varepsilon)))$. In [LS25], the authors showed for any rate $R$, random linear codes are $(1-R-\varepsilon,\ell,L)$-list recoverable, where $L\leq\left(\frac{\ell}{\varepsilon}\right)^{O\left(\frac{\ell}{\varepsilon}\right)}$. Very recently, the work of Brakensiek, Chen, Dhar, and Zhang [BCDZ25a] improved the upper bound on the output list size to $\left(\frac{\ell}{R+\varepsilon}\right)^{O(R/\varepsilon)}$. In their concurrent work [BCDZ25b], they were able to transfer this result to explicit constructions of Folded Reed-Solomon codes and univariate Multiplicity codes as well.
@@
Work Radius Output List Size Alphabet Size
[KRSW23] (Theorem 6.7), [Tam24] (Theorem 4.5) $(1-R-\varepsilon)$ $\left(\ell/\varepsilon\right)^{\left(\frac{\ell}{\varepsilon}\right)^{2}\cdot\log\left(\frac{\ell}{\varepsilon}\right)}$ $\ell^{O(1/\varepsilon^{4})}$
[Tam24] $(1-R-\varepsilon)$ $(\ell/\varepsilon)^{O(\log\ell/\varepsilon)}$ $(n\ell/\varepsilon^{2})^{O(\ell/\varepsilon^{2})}$
[ST25] $(1-R-\varepsilon)$ $\exp(\exp((\ell/\varepsilon)\log(\ell/\varepsilon)))$ $\exp(\exp((\ell/\varepsilon)\log(\ell/\varepsilon)))$
[BCDZ25b] $(1-R-\varepsilon)$ $(\ell/(R+\varepsilon))^{O(R/\varepsilon)}$ $(n\ell/\varepsilon^{2})^{O(\ell/\varepsilon^{2})}$
Our Work $(1-R-\varepsilon)$ $(\ell/(R+\varepsilon))^{O(R/\varepsilon)}$ $\exp\bigl((\ell/\varepsilon)^{(\ell/\varepsilon)^{(\ell/\varepsilon)}}\bigr)$
Table 1: Parameters for recent explicit constructions of list-recoverable codes at capacity. The second and fourth entries hold for both Folded Reed-Solomon codes and Univariate Multiplicity codes.
Table˜1 above highlights the parameters achieved by recent explicit constructions for list-recoverable codes at capacity. We also refer to Table 1 in [BCDZ25a] for parameters achieved by recent randomized constructions. Comparing our result to two works: first, to the construction of [KRSW23] (Theorem 6.7) combined with Tamo’s analysis of the output list size of Folded Reed-Solomon codes [Tam24] (Theorem 4.5), and second, to [ST25], we note that our result attains smaller output list sizes at the cost of increased alphabet size. Compared with [BCDZ25b], both results achieve the same output list size bounds, but their alphabet size grows polynomially with the block length, whereas ours is constant.
All results mentioned below follow from Theorem˜1.1 by specializing to the appropriate list recovery variant.
Theorem 1.6 (Informal, see Corollary˜6.11, Corollary˜6.10).
@@
Instantiating the codes in Theorem˜1.6 with the upper bound on output list sizes for random linear codes, from the recent result of [BCDZ25a], we get the following corollary.
Corollary 1.7 (Informal, follows from Theorem˜1.6).
There exist explicit constructions of $(1-R-\varepsilon,\ell,(\ell/(R+\varepsilon))^{O(R/\varepsilon)})$-list recoverable codes with rate $R-2\varepsilon$, and alphabet size at most $\exp\bigl((\ell/\varepsilon)^{(\ell/\varepsilon)^{(\ell/\varepsilon)}}\bigr)$.
The generality of our main result also yields explicit constructions for related notions such as zero-error list recovery and erasure list recovery. Zero-error list recovery is a special case of list recovery in which the decoding radius is zero. Such codes have found applications in the design of data structures for the heavy hitters problem [DW22]. In the case of erasure list recovery, some of the input lists may contain only the blank symbol, and the objective is to minimize the number of codewords that remain consistent with the non-blank input lists.
Theorem 1.8 (Informal, see Corollary˜6.12).
@@
Theorem 1.9 (Informal, see Corollary˜6.13).
@@
Theorem 1.10 (Informal, see Corollary˜6.14).
@@
We also get a modest improvement in the alphabet size of explicit list-decodable codes approaching the Generalized Singleton Bound. Previously, [JMST25] constructed explicit codes with the same parameters, except with alphabet size equal to $2^{{\mathrm{poly}}(L^{L}/\varepsilon)}$.
@@
The proof in [JMST25] is based on induction, and leverages the expander property of the graph used in the construction. In contrast, ours is a proof by contradiction, and relies on the sampling property of the underlying graph, thereby providing an alternative proof of essentially the same result. Although implied by Theorem˜1.1, we include a proof of Theorem˜1.11 as it more transparently illustrates the ideas behind the proof of the former theorem.
@@
Concurrent Work.
Around the time of release of an earlier version of this paper, Brakensiek, Chen, Dhar, and Zhang released two papers [BCDZ25a, BCDZ25b], independently of our work. We provide a brief discussion about the connections between our work and theirs.
    [BCDZ25a] shows an improved upper bound on the output list sizes achieved by random linear codes, in the context of list-recovery. Specifically, they show that for integers $\ell,L$ and $R,\varepsilon\in[0,1]$, a random linear code with alphabet size $\exp((L\log\ell)/\varepsilon)$ is $(1-R-\varepsilon,\ell,L)$-list-recoverable with high probability, where $L$ satisfies
    $L\leq\left(\frac{\ell}{R+\varepsilon}\right)^{O(R/\varepsilon)}.$
    In the current version of our paper, we have utilized this upper bound to show that our explicit constructions attain the same parameters, with the exception of an increased (yet constant) alphabet size.
    Their result was established by using a novel two-way reduction between subspace designable codes and random linear codes, shown in their second paper [BCDZ25b]. The reduction proves that subspace designable codes (which include explicit Folded Reed-Solomon codes and Univariate Multiplicity codes) achieve all LCL properties simultaneously, at the same level of optimality as random linear codes. Our constructions also achieve the same result (see Remark˜1.3).
    We now compare these two results. The instantiation of Folded Reed-Solomon codes and Univariate Multiplicity codes in [BCDZ25b] requires an alphabet size that is at least polynomial in the block length of the code, while our constructions require constant alphabet size. Additionally, our codes are LDPC codes, a property useful in designing encoding and decoding algorithms. However, our constructions have weaker parameters than [BCDZ25b] in the regime of very large alphabet size, which is a regime of interest in their paper. The constructions of [BCDZ25b] require alphabet sizes that can be significantly larger than the block length, in order to obtain results on tensor codes, as detailed in the second part of their paper.
@@ 1.2 Organization
In Section˜2, we provide an overview of the proofs for the list-decoding case (Theorem˜1.11) and the general result (Theorem˜1.1). The detailed proofs appear in Section˜4 and Section˜5, respectively. Section˜5 is independent of Section˜4; therefore, readers interested solely in the general theorem’s proof may proceed directly to it. Finally, the results pertaining to list-recovery variants and perfect hash matrices (Theorem˜1.6, Theorem˜1.8, Theorem˜1.9, Theorem˜1.10) are presented in Section˜6.
@@ 2 Technical Overview
We begin with an overview of the proof of Theorem˜1.11, which addresses the special case of list decoding. We instantiate the AEL procedure with the following three components: an inner code of constant block length, found through brute force, an explicit outer code having high rate and distance $\delta_{\textnormal{out}}$, where $\delta_{\textnormal{out}}$ is a constant, and an explicit bipartite expander graph $G$. We note that $G$ can equivalently be interpreted as a sampler, a viewpoint that has been leveraged in prior works (cf. [KMRS17, KRSW23]) to analyze AEL-based constructions. This sampling perspective naturally motivates a strategy for the explicit construction of codes with strong list-decoding parameters: perform a brute-force search to identify a constant block-length code with good list-decoding parameters, and then use this code as the inner code in the AEL procedure to obtain $\mathcal{C}_{\textnormal{AEL}}$. If $\mathcal{C}_{\textnormal{AEL}}$ has pairwise distinct codewords $c_{1},\ldots,c_{L+1}$ close to some received word $y$, then the codewords agree with $y$ at a large number of coordinates. The nature of these agreements can be encoded by an agreement hypergraph on $L+1$ vertices that contains $n$ hyperedges, one for each coordinate. The hyperedge for a coordinate is simply the set of indices of codewords agreeing with $y$ on that coordinate. It follows that if $c_{1},\ldots,c_{L+1}$ agree with $y$ on many coordinates, then the sum of the sizes of the hyperedges is large. Let these agreements be described by an agreement hypergraph denoted by $\mathcal{H}$. Upon invoking the sampling property of the graph $G$, it is observed that the precise pattern of these agreements is “ported over” to a significant number of vertices on the left. As a result, the projections of codewords $c_{1},\ldots,c_{L+1}$ and $y$ onto several left vertices have agreements that closely resemble the agreement pattern described by $\mathcal{H}$. Consequently, the codeword projections agree with the projection of $y$ at a large number of coordinates within the inner code. But as the codeword projections are also codewords of the inner code, and because the inner code has good list-decoding parameters, its codewords do not have a lot of agreements with the local projection of $y$. Therefore, we have seemingly arrived at a contradiction.
@@
We now search for a linear code of constant block length satisfying the above property, and use it as an inner code. The analysis now proceeds in a manner analogous to the first approach. If $L+1$ codewords from $\mathcal{C}_{\textnormal{AEL}}$ have a large number of agreements with some received word $y$, then the agreement hypergraph contains a weakly-partition-connected hypergraph $\mathcal{H}$. Consequently, a subset of those codewords satisfy the constraints as set forth by $\mathcal{H}$. We then use the sampling property of the graph $G$ to port over information about the nature of the agreements to the left side, with the result that the inner codewords at several vertices on the left satisfy the constraints described by $\mathcal{H}$. From the parameters that we eventually set, it is seen that the number of such parts exceeds $n(1-\delta_{\textnormal{out}})$ (where $\delta_{\textnormal{out}}$ is the minimum distance of the outer code). Consequently, at least one set of inner codewords from such a part must be non-trivial, thereby contradicting our assumption on the inner code.
@@
Every linear constraint set can be written down as a matrix of dimension $L\times L$. Two linear constraint sets $S$ and $T$ are said to be of the same type if there exists a full rank linear transformation that maps the matrix associated with $S$ to that associated with $T$. Even though the number of types can be exponential in $L^{2}$, it is independent of the block length of $\mathcal{C}_{\textnormal{AEL}}$. This fact ensures that if there are $L$ pairwise distinct codewords in $\mathcal{C}_{\textnormal{AEL}}$ that satisfy the linear constraints described by some local profile $M$ associated with $\mathcal{P}$, then upon arranging them in an $n\times L$ matrix, a significant fraction of the coordinates can be partitioned according to the constraint type satisfied by the row whose index is equal to the coordinate. Moreover, the fractional size of each of these sets will be a constant. Thus, the sampling property of the graph $G$ can be utilized to port over the proportion of these types onto several projections on the left. The local projections, which are also codewords belonging to the inner code, consequently have codewords that satisfy (a close approximation of) the local profile $M$. If the AEL procedure were instantiated with an inner code that avoids vector sets satisfying close approximations of the local profiles associated with $\mathcal{P}$, one might hope to derive a contradiction. However, we encounter the same obstacle as before: the guarantee for the inner code applies only when the inner codewords are pairwise distinct, and we lack a mechanism to ensure that this condition is met.
@@
Recall that we are interested in codes that satisfy the complement of an LCL property $\mathcal{P}$. We shall prove in Section˜5 that if random linear codes of rate $R$ satisfy the complement of $\mathcal{P}$, then random linear codes of rate $R$ also satisfy the following property: with high probability, they do not contain any non-zero matrices satisfying any implied local profile $I(M)$, where $M$ is any local profile associated with $\mathcal{P}$. Upon instantiating the AEL procedure with an inner code satisfying the aforementioned property, the analysis can be carried out in a manner analogous to that of the naive approach detailed above. If $L$ pairwise distinct codewords in $\mathcal{C}_{\textnormal{AEL}}$ satisfy a local profile $M$ associated with $\mathcal{P}$, then upon arranging these codewords in a $n\times L$ matrix and applying the appropriate linear map to each of its rows, the matrix obtained satisfies the constraints set forth by the implied local profile $I(M)$. Upon porting the constraint types to the left as before, we see that the projected matrices on to several left vertices are codewords belonging to the inner code, and moreover, they satisfy I(M)I(M) [^2] In reality, the projected matrices satisfy a close approximation of $I(M)$, instead of $I(M)$. Nevertheless, the inner codes we employ are chosen to be sufficiently robust to accommodate this technical subtlety.. The parameters are instantiated in a way so as to ensure that this occurs on more than $(1-\delta_{\textnormal{out}})n$ left vertices, and therefore there is one projected matrix that is non-zero. But because the columns of the projected matrices are codewords of the inner code, and because the inner code avoids containing all non-zero matrices satisfying $I(M)$ for any $M$ associated with the LCL property $\mathcal{P}$, we successfully arrive at a contradiction.